# CyberHost: Taming Audio-driven Avatar Diffusion Model with Region Codebook Attention

### v1

@@ 1 Introduction
For the first challenge, most existing works (Zhang et al., 2024; Huang et al., 2024) focus on enhancing the ability of diffusion models to generate key local regions, such as hands and faces, in a video-driven setting. They utilize skeleton priors as conditional inputs to simplify the modeling of body shape and motion patterns, allowing the model to focus primarily on reconstructing body textures. However, video-driven body animation typically requires motion generation modules to create pose templates, and employ retargeting technologies to address skeletal discrepancies, making it inconvenient to use. Additionally, most optimization techniques tailored for video-driven settings cannot be seamlessly adapted to audio-driven settings. Regarding the second challenge, a few studies (Liao et al., 2020; Wang et al., 2023b; Hogue et al., 2024; Corona et al., 2024) have attempt to reduce uncertainty by explicitly establishing the connection between audio signals and motion through a two-stage system. It consists of an audio-to-pose module and a pose-to-video module, using poses or meshes as intermediate representations. Nevertheless, this approach faces several critical limitations: (1) The two-stage framework design increases system complexity and reduces the model’s learning efficiency. (2) The poses or meshes carry limited information related to expressiveness, constraining the model’s ability to capture subtle human nuances. (3) Potential inaccuracies in pose or mesh annotations can diminish the model’s performance. Therefore, there is an urgent need to explore how to optimize the generation quality of key human regions within a one-stage audio-driven framework, while also reducing instability issues caused by the motion uncertainty.
@@ 2 Related Work
Video Generation. Benefiting from the advancements in diffusion models, video generation has made significant progress. Some works (Singer et al., 2022; Blattmann et al., 2023a; Zhou et al., 2022; He et al., 2022; Wang et al., 2023a) have extended the 2D U-Net pretrained on image-to-text task to 3D, achieving temporal video generation. AnimateDiff (Guo et al., 2023) trained a pluggable temporal module on large-scale video data, which can be integrated into other text-to-image diffusion models fine-tuned on specific datasets to achieve temporally smooth video generation. For controllability, VideoComposer (Wang et al., 2024c) trained a composer fusion encoder used to combine flexible conditions. Compared with UNet-based methods, DiT-based approaches have shifted the model architecture from convolutional networks to transformers. Some works, such as EasyAnimate (Xu et al., 2024a), CogVideX (Yang et al., 2024), and Sora (Tim et al., 2024), have expanded the 2D DiT framework to 3D for video generation by incorporating a specialized motion module block. Additionally, to efficiently model video data, some works have designed 3D Variational Autoencoders (VAEs) (Kingma & Welling, 2013) to compress videos along both spatial and temporal dimensions. To accelerate video generation, some methods (Podell et al., 2023; Ren et al., 2024) have proposed lightweight distillation schemes, achieving real-time performance in specific domains.
@@
Body Animation. Existing body animation approaches mainly  (Hu et al., 2023; Wang et al., 2024b; Xu et al., 2023; Karras et al., 2023; Zhou et al., 2022) focus on video-driven settings, where the driving signal is a video sequence. DreamPose (Karras et al., 2023) uses DensePose (Güler et al., 2018) as a control signal and trains a diffusion model to perform pose transfer for any given image, thereby generating human video frames sequentially. MagicAnimate (Xu et al., 2023) extends a 2D U-Net to 3D and fine-tunes it on human body data, thereby enhancing the temporal smoothness of human video generation. AnimateAnyone (Hu et al., 2023) uses skeleton maps as control signals for its diffusion model and employs a dual U-Net architecture to maintain consistency between the generated video and the reference images. Some speech-driven body animation works, GAN-based or Diffusion-based  (Liao et al., 2020; Ginosar et al., 2019; Wang et al., 2023b; Corona et al., 2024) do exist, but these methods are typically two-stage processes. Speech2Gesture (Ginosar et al., 2019) first predicts body gesture sequence and then utilizes a pre-trained GAN to render it to final video. Vlogger (Corona et al., 2024) proposed two diffusion models for human-to-3d-motion and motion-to-2d-video. Two-stage methods require explicit representations as intermediate variables. While the use of intermediate representations can reduce the complexity of training, these representations are often not detailed enough, and the introduced errors lower the performance ceiling. Unfortunately, end-to-end approaches have not been widely explored so far.

### v2

@@ 1 Introduction
For the first challenge, most existing works (Zhang et al., 2024; Huang et al., 2024) focus on enhancing the ability of diffusion models to generate critical human body parts in a video-driven setting. They utilize skeleton priors as conditional inputs to simplify the learning of local structures and motion patterns, allowing the model to primarily focus on the texture reconstruction of key regions. However, video-driven body animation typically requires motion generation modules to create pose templates, and employ retargeting technologies to address skeletal discrepancies, making it inconvenient to use. Additionally, most optimization techniques tailored for video-driven settings cannot be seamlessly adapted to audio-driven settings. Regarding the second challenge, a few studies (Liao et al., 2020; Wang et al., 2023b; Hogue et al., 2024; Corona et al., 2024) have attempt to reduce uncertainty by explicitly establishing the connection between audio signals and motion through a two-stage system. It consists of an audio-to-pose module and a pose-to-video module, using poses or meshes as intermediate representations. Nevertheless, this approach faces several critical limitations: (1) The two-stage framework design increases system complexity and reduces the model’s learning efficiency. (2) The poses or meshes carry limited information related to expressiveness, constraining the model’s ability to capture subtle human nuances. (3) Potential inaccuracies in pose or mesh annotations can diminish the model’s performance. Therefore, there is an urgent need to explore how to optimize the generation quality of critical human body parts within a one-stage audio-driven framework, while also reducing instability issues caused by the motion uncertainty.
@@ 2 Related Work
Video Generation. Benefiting from the advancements in diffusion models, video generation has made significant progress in recent years. Some early works (Singer et al., 2022; Blattmann et al., 2023a; Zhou et al., 2022; He et al., 2022; Wang et al., 2023a) have attempted to directly extend the 2D U-Net pretrained on text-to-image tasks into 3D to generate continuous video segments. AnimateDiff (Guo et al., 2023) trained a pluggable temporal module on large-scale video data, allowing easy application to other text-to-image backbones and enabling text-to-video generation with minimal fine-tuning. For controllability, VideoComposer (Wang et al., 2024c) trained a Composer Fusion Encoder to integrate multiple modalities of input as control conditions, thereby making video generation for complex scenes such as human bodies more controllable. Compared with UNet-based methods, DiT-based methods have shown greater potential in video generation task. Some works, such as EasyAnimate (Xu et al., 2024a), CogVideX (Yang et al., 2024), and Sora (Tim et al., 2024), have expanded the 2D DiT framework to 3D for video generation by incorporating a specialized motion module block. Additionally, to alleviate the computational challenges posed by video generation tasks, some works have designed 3D Variational Autoencoders (VAEs) (Kingma & Welling, 2013) for more extreme compression of video information.
@@
Body Animation. Existing body animation approaches mainly  (Hu et al., 2023; Wang et al., 2024b; Xu et al., 2023; Karras et al., 2023; Zhou et al., 2022) focus on video-driven settings, where control signals are the pose sequence extract from the driving video. DreamPose (Karras et al., 2023) uses DensePose (Güler et al., 2018) as a control signal and trains a diffusion model to perform pose transfer for any given image, thereby generating human video frames sequentially. MagicAnimate (Xu et al., 2023) extends a 2D U-Net to 3D and fine-tunes it on human body data, thereby enhancing the temporal smoothness of human video generation. AnimateAnyone (Hu et al., 2023) uses skeleton maps as control signals for its diffusion model and employs a dual U-Net architecture to maintain consistency between the generated video and the reference images. Some speech-driven body animation works, GAN-based or Diffusion-based  (Liao et al., 2020; Ginosar et al., 2019; Wang et al., 2023b; Corona et al., 2024) do exist, but they typically employ a two-stage framework. Speech2Gesture (Ginosar et al., 2019) first predicts body gesture sequence and then utilizes a pre-trained GAN to render it to final video. Similarly, Vlogger (Corona et al., 2024) employs two diffusion models to separately perform audio-to-mesh and mesh-to-video mapping. Two-stage methods require explicit representations as intermediate variables. Although intermediate representations can reduce the training difficulty of audio-driven human video generation models, their limited expressive capabilities can also lower the performance ceiling of the entire system. Unfortunately, the end-to-end training of a one-stage diffusion model for audio-driven body video generation tasks has not been explored yet.